OpenStudy (anonymous):

Match each equation with its solution. Match Term Definition x divided by zero point one equals ten A) x equals one hundred zero point one x equals ten B) x equals zero point one ten x equals one C) x equals ten x divided by one hundred equals zero point one D) x equals one

OpenStudy (maxdrizner):

i can try and help you

OpenStudy (anonymous):

ok @justasktheasianguy

OpenStudy (anonymous):

ok this one is too asian for me to do it so i can't do it

OpenStudy (maxdrizner):

hmm

OpenStudy (maxdrizner):

try a website called mathway it should help you out

OpenStudy (maxdrizner):

sorry if i couldn't help you out

OpenStudy (anonymous):

Which inequality is represented by this graph? A number line shows the numbers zero to ten. There is a closed dot at four, and an arrow extending from that closed dot toward the right. two x is greater than or equal to eight x divided by two is greater than or equal to four three x is less than twelve x divided by four is less than or equal to one

OpenStudy (texaschic101):

x/0.1 = 10 --- multiply by 0.1 x = 10(0.1) x = 1(D) 0.1x = 10 --- divide by 0.1 x = 10/0.1 x = 100 (A) 10x = 1 -- divide by 10 x = 1/10 x = 0.1 (B) x/100 = 0.1 -- multiply by 100 x = (0.1)(100) x = 10 (C) =============== closed dot at 4 -- means there is an equal sign in the problem shading to the right -- means it is greater then your solution is : x >= 4 find the answer choice that gives you this answer... 2x >= 8 -- divide by 2 x >= 4 answer is : 2x >= 8

OpenStudy (anonymous):

D A B C :)